# High-Throughput Quantification of Altermagnetic Band Splitting

### v1

@@ III.2 NbMnP
NbMnP crystallizes in the TiNiSi-type orthorhombic structure belonging to the paramagnetic space group Pnma (No. 62), characterized by lattice parameters $a=6.182\,\mathrm{\AA }$, $b=3.557\,\mathrm{\AA }$, and $c=7.219\,\mathrm{\AA }$ [37]. Its crystal structure comprises zigzag chains of Mn atoms aligned along the $b$-axis, embedded within a network formed by Nb and P atoms. In this study, we consider a collinear AFM configuration with Mn magnetic moments oriented along the $z$-axis, as depicted in Fig.3(a).
@@ III.3 YRuO3
IV Comparison with other studies
The results reported here can be discussed in relation to the high–throughput study by Wan et al. [25], which identified 336 potential altermagnets by screening MAGNDATA. Their classification relied on the absence of the antiunitary symmetries $P\!\cdot T$ (inversion plus time reversal) and $t\!\cdot T$ (translation plus time reversal). Moreover, they extended their candidate pool to include ferrimagnetic systems by considering their SOC-off configuration. Our study employs the amcheck algorithm, which tests the same symmetry condition (absence of $P\!\cdot T$ and $t\!\cdot T$) to identify potential altermagnetic materials. However, amcheck strictly excludes ferrimagnetic states since their sublattices are completely not symmetry-related and are instead classified as “Luttinger ferrimagnets” that exhibit spin splitting but lack the required momentum-odd, nodal spin texture over the Brillouin zone. For example, Cr2S3, listed as “interesting” in Ref. [25], shows spin splitting in the spin–polarized DFT consistent with ferrimagnetic behavior, yet lacks the nodal spin texture and is therefore not classified as altermagnetic here. On the other hand, we find a collinear AFM with Kramers–degenerate bands for SrMnSb2, and the small splittings of a few meV that appear at high energy are most likely numerical artifacts and not indicative of altermagnetism. These distinctions likely account for the smaller number of altermagnets identified in our study compared to Ref. [25], while remaining fully consistent with the symmetry–based definition of altermagnetism [1, 2].
@@ IV Comparison with other studies
Table 1: Top altermagnet candidate materials with respect to their spin-splitting. Space group, spin-splitting metrics, band gap, Hubbard $U$, and minimum elemental abundance.
@@
Material Space Group $F_{\Delta}$ (%) Max. SS (eV) Avg. SS (eV) Bandgap (eV) $U$ Values Min Abund.
@@
V Conclusion
In this work, we have established a comprehensive and symmetry-aware high-throughput framework for the discovery and characterization of altermagnetic materials. We systematically screened more than 2,000 experimentally verified magnetic structures from the MAGNDATA database using a two-stage workflow. This approach, built on the magnetic symmetry diagnostic tool amcheck and the High-Throughput Toolkit, identified 171 candidate materials exhibiting significant momentum-dependent spin splitting. These include both metallic and semiconducting systems, many of which had not been previously reported as altermagnets. Notably, our approach recovers experimentally confirmed compounds such as CrSb, MnTe, and RuO2, and highlights previously overlooked materials like UCr2Si2C, MnNbP, and YRuO3 that display robust symmetry-protected spin splitting even in the absence of spin–orbit coupling.
@@ Appendix A Computational Details
Density functional theory (DFT) calculations were carried out using the Vienna Ab initio Simulation Package (VASP) [33, 32], employing the projector augmented-wave (PAW) method and the generalized gradient approximation (GGA) with the Perdew–Burke–Ernzerhof (PBE) functional to describe exchange-correlation effects [31, 39]. Electron correlation was treated using the rotationally invariant DFT+$U$ approach [40], with the effective Hubbard $U_{\text{eff}}$ parameters taken from Pymatgen (Materials Project) [41] and supplemented by Wang et al. [42] where needed.
@@
High-throughput calculations were managed using httk [30]. High-symmetry $\mathbf{k}$-paths were constructed using Pymatgen [29], and magnetic space groups were identified using the Spglib library. Fermi surfaces and spin textures were visualized using the IFermi package [43].

### v2

@@ III.2 NbMnP
NbMnP crystallizes in the TiNiSi-type orthorhombic structure belonging to the paramagnetic space group Pnma (No. 62), characterized by lattice parameters $a=6.182\,\mathrm{\AA }$, $b=3.557\,\mathrm{\AA }$, and $c=7.219\,\mathrm{\AA }$ [41]. Its crystal structure comprises zigzag chains of Mn atoms aligned along the $b$-axis, embedded within a network formed by Nb and P atoms. In this study, we consider a collinear AFM configuration as depicted in Fig. 3(a).
@@ III.3 YRuO3
IV Identification in 2D databases
To identify bulk altermagnetic compounds that may have chemically equivalent two-dimensional counterparts, we compared the reduced chemical formulas of all 188 bulk AM candidates with those reported in two major 2D materials databases: the Computational 2D Materials Database (C2DB) [43] and the AiiDA 2D repository [44]. In Table 1, we collect all bulk AM materials whose reduced formula appears in at least one of the 2D databases and indicate whether a match was found in C2DB, in AiiDA, or in both. These databases contain both experimentally synthesized and theoretically predicted monolayer compounds.
Table 1: Bulk altermagnetic materials with an existing 2D compound of the same reduced chemical formula in the C2DB or AiiDA databases. Multiple MCIF entries with identical formulas correspond to distinct MAGNDATA models.
IDs Chemical formula C2DB AiiDA
0.15 MnF2 Yes
0.178, 0.334 CoF2 Yes
0.335, 0.581, 1.0.33 FeF3 Yes
0.36 NiF2 Yes
0.607 RuO2 Yes
0.800 MnTe Yes
1.0.47, 1.0.48 MnSe2 Yes Yes
2.35 CrSe Yes
0.65, 0.66 Fe2O3 Yes
The list of bulk altermagnets with a corresponding two-dimensional compound is rather short. Indeed, the majority of compounds appearing in Table 1 contain transition–metal elements with partially filled $d$ shells together with halogens (F, Cl) or chalcogens (S, Se, Te). The later help stabilizing layered or quasi-layered chemistries, while the former provide the local moments that form the basis of the magnetism.
Note that the first column of the table shows that some chemical formulas correspond to multiple MAGNDATA entries. The reason is that different experimental studies report distinct magnetic configurations, magnetic subgroup assignments, or slightly different crystallographic refinements. These duplicated entries correspond to separate magnetic models of the same material, not different compounds, and therefore belong match with a single two-dimensional material.
Table 1 identifies candidate bulk altermagnets which exist in 2D, but does not say that these materials are also altermagnetic as a monolayer. This question has been addressed by Sødequist and Olsen [45], who identified 7 altermagnetic monolayer candidates starting from C2DB. Crucially, they take into account the magnetic anisotropy coming from spin-orbit that is necessary to get a magnetically ordered phase in 2d, in accordance with the Mermin-Wagner theorem. Starting from C2DB, they identified which monolayers can be altermagnets, also taking into account spin-orbit coupling. The spin-orbit coupling introduces magnetic anisotropy that makes it possible to stabilize magnetism in 2d while satisfying the Mermin-Wagner theorem. Their 7 altermagnetic monolayer candidates similarly contain mostly transition metals terminated by halogens or chalcogens, although they also find two more complex compounds. Interestingly, there is zero overlap between Table 1 and the list in Sødequist and Olsen.
V Comparison with other studies
[…]
Furthermore, we have considered a set of collinear versions of materials reported as noncollinear. This extended set of materials along with their computed spin splittings describe alternative spin configurations on otherwise stable crystal structures. They can be a valuable resource to aid the general understanding of altermagnetism. Some of them are arguably relevant due to them being very nearly collinear, shown in our analysis by the angle between the spin axes. We also speculate that for some of the ones with larger such angles, there could exist ways in practice to invoke a more aligned moments to give the altermagnetic configuration, via, e.g., external fields, stress/strain, or similar mechanisms.
@@ V Comparison with other studies
Table 2: Top altermagnet candidate materials with respect to their spin-splitting. Space group, max angle between spins for materials that were originally noncollinear ($\arccos(|S_{i}\cdot S_{j}|/|S_{i}||S_{j}|)$), spin-splitting metrics, band gap, wave classification, Hubbard $U$, and minimum elemental abundance. For more information on symmetry data, refer to the supplementary materials.
@@
Material Space Group $\frac{|S_{i}\cdot S_{j}|}{|S_{i}||S_{j}|}$ $F_{\Delta}$ (%) Max. SS (eV) Avg. SS (eV) Bandgap (eV) Wave $U$ Values Min Abund.
@@
VI Conclusion
In this work, we have established a comprehensive and symmetry-aware high-throughput framework for the discovery and characterization of altermagnetic materials. We systematically screened more than 2,000 experimentally verified magnetic structures from the MAGNDATA database using a two-stage workflow. This approach, built on the magnetic symmetry diagnostic tool amcheck and the High-Throughput Toolkit, identified 171 candidate materials exhibiting significant momentum-dependent spin splitting. These include both metallic and semiconducting systems, many of which had not been previously reported as altermagnets. Notably, our approach recovers experimentally confirmed compounds such as CrSb, MnTe, and RuO2, and highlights previously overlooked materials like UCr2Si2C, MnNbP, and YRuO3 that display robust symmetry-protected spin splitting even in the absence of spin–orbit coupling. Furthermore, by cross-referencing our bulk AM candidates with the C2DB and AiiDA 2D repository, we identify nine materials with chemically equivalent 2D counterparts.
@@ Appendix A Computational Details
Density functional theory (DFT) calculations were carried out using the Vienna Ab initio Simulation Package (VASP) [37, 36], employing the projector augmented-wave (PAW) method and the generalized gradient approximation (GGA) with the Perdew–Burke–Ernzerhof (PBE) functional to describe exchange-correlation effects [35, 46]. Electron correlation was treated using the rotationally invariant DFT+$U$ approach [47], with the effective Hubbard $U_{\text{eff}}$ parameters taken from Pymatgen (Materials Project) [48] and supplemented by Wang et al. [49] where needed. The magnitude of the spin splitting depends on the value of $U$ via the magnitude of the magnetic moments [50], whereas the structure of the splitting in the Brillouin Zone is expected to be weakly dependent on the value of $U$.
@@
High-throughput calculations were managed using httk [34]. High-symmetry $\mathbf{k}$-paths were constructed using Pymatgen [33], and magnetic space groups were identified using the Spglib library [31, 32]. Fermi surfaces and spin textures were visualized using the IFermi package [51].